OpenStudy (anonymous):

A taxicab starts at (1, -2) on the grid and picks up a passenger. The taxicab travels a total of 16 blocks, and then drops the passenger off. Which composition of two translations could describe the taxicab’s final position?

OpenStudy (jdoe0001):

is there a picture of "the grid", and which choices are there?

OpenStudy (anonymous):

(x, y) (x - 4, y + 3), and then (x, y) (x + 6, y - 1). (x, y) (x - 3, y + 4), and then (x, y) (x + 7, y - 1). (x, y) (x + 4, y - 3), and then (x, y) (x - 9, y + 1). (x, y) (x - 3, y + 4), and then (x, y) (x + 8, y - 1).

OpenStudy (jdoe0001):

https: :/

OpenStudy (anonymous):

thats the URL

OpenStudy (jdoe0001):

the url works, usually they don't

OpenStudy (jdoe0001):

doesn't shed much light on the blocks though

OpenStudy (anonymous):

thats all my math book gives me :/

OpenStudy (jdoe0001):

well, if I understand correctly, the "moves" or "offsets" are from the "current position" so, let see the 1st one (x, y) (x - 4, y + 3), it went "4 units to the left" over the x-axis, then "3 units to the right" over the y-axis THEN (x, y) (x + 6, y - 1). it went from there, "6 units to the right" over x-axis,, and "1 unit to the left" over the y-axis so the taxicab went, 4+6 units over the x-axis 3+1 units over the y-axis so, 4+6+3+1 \(\ne\) 16 so, that's not it

OpenStudy (jdoe0001):

let's see the 2nd one (x, y) (x - 3, y + 4), and then (x, y) (x + 7, y - 1). 3+7+4+1 \(\ne\) 16 so it's not that one either

OpenStudy (jdoe0001):

the 3rd option (x, y) (x + 4, y - 3), and then (x, y) (x - 9, y + 1). 4+9+3+1 \(\ne\) 16 no dice there too

OpenStudy (jdoe0001):

last but not least (x, y) (x - 3, y + 4), and then (x, y) (x + 8, y - 1). 3+8+4+1 = 11+5 = 16

OpenStudy (anonymous):

thank you so much that problem has been bugging me that i have not been able to solve it

OpenStudy (jdoe0001):

yw